# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v2

@@ 5.3.2 Operator-perspective results
Table 4 shows the operator-perspective results alongside results from other work. Kılıç and Gök (2014) do not report results for the operator perspective and as such we do not include their work in Table 4. Similarly to the passenger-perspective results, our methods underperform on the smallest cities (Mandl and Mumford0) but perform well on larger ones. Strikingly, NEA achieves the lowest value of total route time $C_{o}$ out of all methods on Mumford1, 2, and 3, improving even on Hüsselmann et al. (2023)’s NSGA-II. Evidently, the learned policy functions very well as a low-level heuristic at low values of $\alpha$, where the premium is on keeping routes short.
@@
The metrics $d_{0}$ to $d_{un}$ reveal that both PC-EA and NEA favour higher numbers of transfers relative to most of the other methods, particularly Hüsselmann et al. (2023). This is true for both the passenger and operator perspectives. For the operator perspective, this matches our intuitions: shorter routes will deliver fewer passengers directly to their destinations. But we were surprised to observe this in the passenger perspective case as well.
Compared to Hüsselmann et al. (2023)’s DBMOSA, PC-EA achieves lower average trip time $C_{p}$ on Mumford2 and Mumford3 for the passenger perspective, despite the fact that it requires more passenger trips to make 1 and 2 transfers. Still, PC-EA does give fewer transfers overall than NEA, which matches our intuition that fewer transfers should result in lower average trip time $C_{p}$. It seems that the process of compositing shortest paths to form routes biases those routes towards directness and efficiency, to a degree that outweighs the cost imposed by having more transfers. Each transfer is less impactful on riders if the routes they’re transferring between are more direct.
Meanwhile, since both the learned heuristic of NEA and the unlearned heuristic of PC-EA suffer from high transfer counts, it seems the difference between these and the other results from the literature may be a result of the evolutionary algorithm itself. At any rate, these results are not out of line with those of Mumford (2013a) and Kılıç and Gök (2014) on the largest cities, so we do not consider this a major strike against our learned heuristics. Indeed, we expect that these results could be improved by explicitly including $d_{0}$ in the reward function used to train the policy $\pi_{\theta}$, and this would reduce the number of transfers required by NEA’s networks.
@@ 6.3 Experiments
We performed experiments evaluating both the NEA and PC-EA algorithms on the Laval city graph. Three sets of experiments were performed with different $\alpha$ values representing different perspectives: the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). As in section 5, we perform ten runs of each experiment with ten different random seeds and ten different learned policies $\pi_{\theta}$, and report statistics over these ten runs; the same per-seed parameters $\theta$ are used as in the experiments of section 5. We also run the STL network on the Laval city graph $\mathcal{C}$ in the same way that we run the networks from our other algorithms to measure their performance, but since the STL network is a single, pre-determined network that does not depend on $\alpha$, we run it only once.
Some trial runs we performed with $IT=4,000$ showed that the network cost stopped decreasing after $2,700$ iterations, so to save time we ran our final NEA and PC-EA experiments on Laval with $IT=3,000$. The other parameters of the algorithm were the same as in the Mumford experiments: $B=10$, $E=10$, and transfer penalty $p_{T}=300$s.
@@ 6.3.1 Results
Table 6: Performance of networks from LC-100, NEA, and PC-EA at three $\alpha$ values, as well as the STL network. Bolded values are best for the corresponding $\alpha$. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satsified with 3 or more transfers. All values are averaged over ten random seeds, with one standard deviation following “$\pm$”, except for those in the STL row.
@@
Figure 7: Average trip time $C_{p}$ versus total route time $C_{o}$ achieved by LC-100, NEA, and PC-EA, for $\alpha=0.0$ (bottom-right), 0.5 (middle), and 1.0 (top-left), as well as for the STL network itself (independent of $\alpha$). Except for STL, each point is a mean value over 10 random seeds, and bars show one standard deviation. Note that PC-EA’s $\alpha=0.0$ and 0.5 have nearly identical values (as shown in Table 6), so the corresponding points are stacked in this figure and appear as one point.
@@
We see that for each value of $\alpha$, NEA’s network outperforms the STL network: it achieves 54% lower $C_{p}$ at $\alpha=1.0$, 18% lower $C_{o}$ at $\alpha=0.0$, and at $\alpha=0.5$, 34% lower $C_{p}$ and 12% lower $C_{o}$. At both $\alpha=0.0$ and $\alpha=0.5$, NEA strictly dominates both lower $C_{p}$ and lower $C_{o}$ than the existing transit system.
@@
Comparing NEA’s results to LC-100, we see that at $\alpha=0.0$, NEA decreases $C_{o}$ and increases $C_{p}$ relative to LC-100; and for $\alpha=1.0$, the reverse is true. In both cases, NEA improves over the initial network on the objective being optimized, at the cost of the other, as we would expect. In the balanced case ($\alpha=0.5$), we see that NEA decreases $C_{o}$ by 13% while increasing $C_{p}$ by 4%. This matches what we observed in Figure 3, as discussed in subsection 5.1.
@@
Looking at the transfer percentages $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$, we see that as $\alpha$ increases, the overall number of transfers shrinks, with $d_{1}$ growing and $d_{un}$ shrinking. This is as expected, since fewer transfers usually means shorter passenger travel times. By the same token, PC-EA performs best on these metrics, as it does on average trip time $C_{p}$. Notably though, at $\alpha=0.0$ NEA’s $d_{0}$, $d_{1}$, and $d_{2}$ are only slightly worse than the STL network’s - the largest increase is in $d_{2}$, by 13% of STL’s value - while $d_{un}$ is less than the STL network’s, meaning that NEA’s output networks let more passengers reach their destination in 2 or fewer transfers than the STL network. In fact, the very small decrease in $d_{0}$ implies that the larger increases in $d_{1}$ and $d_{2}$ mostly come from trips that were in $d_{un}$ in the STL network; NEA has thus decreased the average number of transfers. This is a significant improvement, given that at $\alpha=0.0$ the algorithm is unconcerned with minimizing passenger travel time.
At $\alpha=0.5$, NEA’s $d_{un}$ decreases further and $d_{0}$, $d_{1}$, and $d_{2}$ correspondingly increase further. Then from $\alpha=0.5$ to $1.0$, NEA’s $d_{un}$ drops by more than half, and its $d_{0}$, $d_{1}$, and $d_{2}$ are all markedly increased, especially $d_{0}$ which is 31% greater than STL’s $d_{0}$. NEA improves on the STL network by a considerable margin at all three $\alpha$ values, even though PC-EA outperforms NEA at $\alpha=1.0$.
@@
The likely aim of Laval’s transit agency is to reduce costs while improving service quality, or at least not harming it too much. The “balanced” case, where $\alpha=0.5$, is most relevant to that aim. In this balanced case, we find that NEA proposes networks that reduce total route time by 13% versus the existing STL transit network. Some simple calculations can help understand the savings this represents. The approximate cost of operating one of Laval’s buses is 200 Canadian dollars (CAD) per hour. Suppose that each route has the same headway (time between bus arrivals) $H$. Then, the number of buses $N_{r}$ required by route $r$ is the ceiling of the route’s driving time $\tau_{r}$ divided by the headway, $\left\lceil\frac{\tau_{r}}{H}\right\rceil$. The total number of buses $N_{\mathcal{R}}$ required for a transit network $\mathcal{R}$ is then:
@@
The total cost of the system is 200 CAD×$\times N_{\mathcal{R}}$. Assuming a headway of 15 minutes on all routes, that gives a per-hour operating cost of 319,387 CAD for the STL network, while the average operating cost of networks from NEA with $\alpha=0.5$ is 279,707 CAD. So, under these simplifying assumptions, NEA could save the Laval transit agency on the order of 40,000 CAD per hour - savings of more than 12% of costs. This is only a rough estimate: in practice, headways differ between routes, and there may be additional costs involved in altering the routes from the current system, such as from moving or building bus stops. But it shows that NEA, and neural heuristics more broadly, may offer practical savings to transit agencies.
In addition, NEA’s balanced-case networks reduce the average passenger trip time by 34% versus the STL network, and reduce the number of transfers that passengers have to make, especially reducing the percentage of trips that require three or more transfers from 41.83% to 36.41%. Trips requiring three or more transfers are widely regarded as being so unattractive that riders will not make them (but note that these trips are still included in the calculation of $C_{p}$). So NEA’s networks may even increase overall transit ridership. This not only makes the system more useful to the city’s residents, but also increases the agency’s revenue from fares.
We also note that PC-EA achieves lower average trip time $C_{p}$ than both STL and NEA for all three $\alpha$ values, but at the cost of worsening total route time $C_{o}$ versus STL in each case. Furthermore, at $\alpha=1.0$ where we disregard $C_{o}$, it achieves only 3% lower $C_{p}$ than NEA. This supports our conclusions in subsection 5.2.3 and subsection 5.3 that the learned heuristics of the neural net policies $\pi_{\theta}$ are the main driver of NEA’s performance, rather than simply the process of assembling routes from shortest paths.
In these experiments, the nodes of the graph represent census dissemination areas instead of existing bus stops. In order to be used in reality, the proposed routes would need to be fitted to the existing locations of bus stops in the CDAs, likely making multiple stops within each CDAs. To minimize the cost of the network redesign, it would be necessary to keep stops at most or all stop locations that have shelters, as constructing or moving these may cost tens of thousands of dollars, while stops with only a signpost at their locations can be moved at the cost of only a few thousand dollars. An algorithm for translating our census-level routes to real-stop-level routes, in a way that minimizes cost, is an important next step for making this work applicable. We leave this for future work.
@@ 7 CONCLUSIONS
A remaining question is why the random path-combining heuristic used in the PC-EA experiments outperforms the learned heuristic in the extreme of the passenger perspective. As noted in subsection 5.2.3, in principle the neural net should be able to learn whatever policy the random path-combiner is enacting: in the limit, it could learn to give the same probability to all actions when $\alpha=1.0$. We wish to explore in more depth why this does not occur, and whether changes to the policy architecture or learning algorithm might allow the learned heuristic’s performance to match or exceed that of the random policy.
@@ A.2.2 Halting module
$\textup{NN}_{halt}$ is a simple feed-forward MLP with 2 hidden layers. It takes as input the concatenation of the node embeddings of the first and last nodes on the current route, $\mathbf{s}_{t}$ and $\tau_{r_{t}}$. It outputs a scalar $h$. We apply the sigmoid function to $h$ get the halting policy:
@@ A.2.3 Extension module
The extension module $\textup{NN}_{ext}$ is finely tuned to the structure of the cost function. Consider first the case where $\alpha=1$. The benefit of a candidate route $r$ is then solely a function of the edges $(i,j,\tau_{rij})$ that it adds to the set $\mathcal{E}_{\mathcal{R}}$ of direct-transit-connection edges, where $\tau_{rij}$ denotes the time to get from $i$ to $j$ on route $r$. Therefore, for each path $a\in\textup{SP}$ that we consider as an extension of $r$, a reasonable heuristic is to compute a scalar score $o_{ij}$ for each edge $(i,j,\tau_{aij})$, and let the path’s score $o_{a}$ be the sum of the scores of all pairs of nodes that are connected by that path.
Following this intuition, for each node pair $(i,j)$, we concatenate $[\mathbf{y}_{i},\mathbf{y}_{j},\mathbf{e}_{ij},\tau_{(r+a)ij}]$ into a vector and apply an MLP to to compute the scalar $o_{ij}$. For each path $a\in\textup{SP}$, we sum these node-pair scores to get path scores: $o_{a}=\sum_{i,j\in a}o_{ij}$.
@@
If $\alpha<1$, then the quality of $r$ still depends on the edges it adds to $\mathcal{E}_{\mathcal{R}}$, but also on its driving time, and whether it is a good choice to extend $r$ by $a$ depends generally on the state $\mathcal{S}$. So, we concatenate $[o_{a},\tau_{a},\mathbf{s}_{t}]$ and apply another MLP with 2 hidden layers to obtain a final score $\hat{o}_{a}$ for each candidate path. The extension policy is then the softmax of these values:
$\pi(a)=\frac{e^{\hat{o}_{a}}}{\sum_{a^{\prime}\in\mathcal{A}}e^{\hat{o}_{a^{\prime}}}}$ (13)
A.3 Baseline Network

### v3

@@ 5.3.3 Operator-perspective results
Table 4 shows the operator-perspective results alongside results from other work. Kılıç and Gök (2014) do not report results for the operator perspective and as such we do not include their work in Table 4. Similarly to the passenger-perspective results, our methods underperform on the smallest cities (Mandl and Mumford0) but perform well on larger ones. NEA sets a new state of the art in total route time $C_{o}$ on the largest graph Mumford3, and outperforms all methods other than Ahmed et al. (2019) on Mumford1 and Mumford2. NEA’s performance gap with Ahmed et al. (2019) narrows consistently over the Mumford cities as they grow in size, until surpassing it on Mumford3, suggesting that its advantage will continue to grow as the city gets larger. Evidently, the learned policy functions very well as a low-level heuristic at low values of $\alpha$, where the premium is on keeping routes short.
@@
5.3.4 Number of transfers
The metrics $d_{0}$ to $d_{un}$ reveal that both RC-EA and NEA favour higher numbers of transfers relative to most of the other methods, particularly Hüsselmann et al. (2023). This is true for both the passenger and operator perspectives. For the operator perspective, this matches our intuitions: shorter routes will deliver fewer passengers directly to their destinations and will require more transfers. But we were surprised to observe this in the passenger perspective case as well.
Compared to Hüsselmann et al. (2023)’s DBMOSA, RC-EA achieves lower average trip time $C_{p}$ on Mumford2 and Mumford3 for the passenger perspective, despite the fact that it requires more passenger trips to make 1 and 2 transfers. Still, RC-EA’s networks require fewer transfers overall than NEA’s, which matches our intuition that fewer transfers should result in lower average trip time $C_{p}$. It seems that the process of compositing shortest paths to form routes biases those routes towards directness and efficiency, to a degree that outweighs the cost imposed by having more transfers. Each transfer is less impactful on riders if the routes they’re transferring between are more direct.
Meanwhile, since both the learned heuristic of NEA and the unlearned heuristic of RC-EA suffer from high transfer counts, it seems the difference between these and the other results from the literature may be a result of the evolutionary algorithm itself. At any rate, these results are not out of line with those of Mumford (2013a) and Kılıç and Gök (2014) on the largest cities, so we do not consider this a major strike against our learned heuristics. Indeed, we expect that these results could be improved by explicitly including $d_{0}$ in the reward function used to train the policy $\pi_{\theta}$, and this would reduce the number of transfers required by NEA’s networks.
@@ 6.1.1 Street graph
The real city of Laval contains 2,618 unique transit stop facilities. Taking each of these to be a node would yield too large a graph for our neural policy to be able to process on the commercial GPU hardware available to us - the GPU memory requirements would simply be too great. It is also likely that such a representation would be more fine-grained than necessary: the demand for travel within the service areas of nearby transit stops is likely to be very similar. For these reasons, we instead made use of a coarsened graph, where the nodes correspond to CDAs. CDAs are designed so as to enclose populations that are roughly homogeneous (among other factors), making them a sensible granularity at which to analyze travel demands. Laval contains 632 distinct CDAs, and a graph size of $n=632$ makes it feasible for us to apply our neural policy with the hardware we have available.
@@ 6.3 Experiments
We performed experiments evaluating both the NEA and RC-EA algorithms on the Laval city graph. Three sets of NEA experiments were performed with different $\alpha$ values representing different perspectives: the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). RC-EA experiments are only run at $\alpha=1.0$ since we observed in subsection 5.2.3 that RC-EA did not perform well at other $\alpha$. As in section 5, we perform ten runs of each experiment with ten different random seeds and ten different learned policies $\pi_{\theta}$, and report statistics over these ten runs; the same per-seed parameters $\theta$ are used as in the experiments of section 5. We also run the STL network on the Laval city graph $\mathcal{C}$ in the same way that we run the networks from our other algorithms to measure their performance, but since the STL network is a single, pre-determined network that does not depend on $\alpha$, we run it only once.
The parameters of NEA were the same as in the Mumford experiments in subsection 5.1: $IT=400$, $B=10$, $E=10$, and transfer penalty $p_{T}=300$s. While the memory requirements of running our method on this very large graph were greater than for the Mumford benchmark datasets, we were still able to run the experiments on the desktop machine used for our Mumford experiments, without any changes to our model or algorithm code.
@@ 6.3.1 Results
Table 6: Performance of networks from LC-100 and NEA at three $\alpha$ values, RC-EA at $\alpha=1.0$, and the STL network. Bolded values are best for the corresponding $\alpha$, except where STL performs better than all methods at that $\alpha$, in which case no values are bolded. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satisfied with 3 or more transfers. All values are averaged over ten random seeds, with one standard deviation following “$\pm$”, except for those in the STL row.
@@
Figure 7: Average trip time $C_{p}$ versus total route time $C_{o}$ achieved by LC-100 and NEA at $\alpha=0.0$ (bottom-right), 0.5 (middle), and 1.0 (top-left), as well as for RC-EA at $\alpha=1.0$ and the STL network itself (independent of $\alpha$). Except for STL, each point is a mean value over 10 random seeds, and bars show one standard deviation.
@@
We see that for each value of $\alpha$, NEA’s network outperforms the STL network: it achieves 52% lower $C_{p}$ at $\alpha=1.0$, 24% lower $C_{o}$ at $\alpha=0.0$, and at $\alpha=0.5$, 33% lower $C_{p}$ and 17% lower $C_{o}$. At both $\alpha=0.0$ and $\alpha=0.5$, NEA strictly dominates the existing transit system, achieving both lower $C_{p}$ and lower $C_{o}$.
@@
Comparing NEA’s results to LC-100, we see that at $\alpha=0.0$, NEA decreases $C_{o}$ and increases $C_{p}$ relative to LC-100; and for $\alpha=1.0$, the reverse is true. In both cases, NEA improves over the initial network on the objective being optimized, at the cost of the other, as we would expect. In the balanced case ($\alpha=0.5$), we see that NEA decreases $C_{o}$ by 7% while increasing $C_{p}$ by 1.6% relative to LC-100. This matches what we observed in Figure 3, as discussed in subsection 5.1. NEA improves on the STL network by a considerable margin at all three $\alpha$ values, even though RC-EA outperforms NEA at $\alpha=1.0$.
@@
Looking at the transfer percentages $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$, we see that as $\alpha$ increases, the overall number of transfers shrinks, with $d_{0}$, $d_{1}$ and $d_{2}$ growing and $d_{un}$ shrinking. This is as expected, since fewer transfers implies reduced passenger travel times. By the same token, at $\alpha=1.0$ RC-EA performs best on these metrics, as it does on average trip time $C_{p}$. At $\alpha=1.0$, NEA outperforms the STL network by these metrics as well. At $\alpha=0.5$, surprisingly, NEA’s $d_{0}$ is lower than at the other $\alpha$ values, and lower than the STL network’s $d_{0}$ by 0.8% of trips. But NEA’s $d_{un}$ is lower than STL’s by 4.8% of trips, while $d_{1}$ and $d_{2}$ are both higher, meaning that NEA’s networks let more passengers reach their destination in 2 or fewer transfers than the STL network. It is only at $\alpha=0.0$ that STL’s network requires fewer transfers overall than NEA’s networks, and this is reasonable given that at $\alpha=0.0$ the goal is strictly to minimize route length regardless of transfers.
@@
The likely aim of Laval’s transit agency is to reduce costs while improving service quality, or at least not harming it too much. The “balanced” case, where $\alpha=0.5$, is most relevant to that aim. In this balanced case, we find that NEA proposes networks that reduce total route time by 17% versus the existing STL transit network. Some simple calculations can help understand the savings this represents. The approximate cost of operating a Laval city bus is 200 Canadian dollars (CAD) per hour. Suppose that each route has the same headway (time between bus arrivals) $H$. Then, the number of buses $N_{r}$ required by route $r$ is the ceiling of the route’s driving time $\tau_{r}$ divided by the headway, $\left\lceil\frac{\tau_{r}}{H}\right\rceil$. The total number of buses $N_{\mathcal{R}}$ required for a transit network $\mathcal{R}$ is then:
@@
The total cost of the system is 200 CAD×$\times N_{\mathcal{R}}$. Assuming a headway of 15 minutes on all routes, that gives a per-hour operating cost of 319,387 CAD for the STL network, while the average operating cost of networks from NEA with $\alpha=0.5$ is 266,373 CAD. So, under these simplifying assumptions, NEA could save the Laval transit agency on the order of 53,000 CAD per hour, or about 17% of operating costs. This is only a rough estimate: in practice, headways differ between routes, and there may be additional costs involved in altering the routes from the current system, such as from moving or building bus stops. But it shows that NEA, and neural heuristics more broadly, may offer practical savings to transit agencies.
In addition, NEA’s balanced-case networks reduce the average passenger trip time by 33% versus the STL network, and reduce the number of transfers that passengers have to make, especially reducing the percentage of trips that require three or more transfers from 41.83% to 37.08%. Trips requiring three or more transfers are widely regarded as being so unattractive that riders will not make them (but note that these trips are still included in the calculation of $C_{p}$). So NEA’s networks may even increase overall transit ridership. This not only makes the system more useful to the city’s residents, but also increases the agency’s revenue from fares.
We also note that at $\alpha=1.0$, RC-EA achieves only 3% lower $C_{p}$ than NEA. This supports our conclusions in subsection 5.2.3 and subsection 5.3 that the learned heuristics of the neural net policies $\pi_{\theta}$ are the main driver of NEA’s performance across $\alpha$, rather than simply the heuristic of assembling routes from shortest paths.
In these experiments, the nodes of the graph represent census dissemination areas instead of existing bus stops. In order to be used in reality, the proposed routes would need to be fitted to the existing locations of bus stops in the CDAs, likely making multiple stops within each CDAs. To minimize the cost of the network redesign, it would be necessary to keep stops at most or all stop locations that have shelters, as constructing or moving these may cost tens of thousands of dollars, while stops with only a signpost at their locations can be moved at the cost of only a few thousand dollars. An algorithm for translating our census-level routes to real-stop-level routes, in a way that minimizes cost, is an important next step if the methods developed here are to be applied in real cities. We leave this for future work.
@@ 7 CONCLUSIONS
A remaining question is why the random path-combining heuristic used in the RC-EA experiments outperforms the learned heuristic in the extreme of the passenger perspective. As noted in subsection 5.2.3, in principle the neural net should be able to learn whatever policy the random path-combiner is enacting: in the limit, it could learn to give the same probability to all actions when $\alpha=1.0$. We wish to explore in more depth why this does not occur, and whether changes to the policy architecture or learning algorithm might allow the learned heuristic’s performance to match or exceed that of the random policy.
@@ A.2.2 Halting module
$\textup{NN}_{halt}$ is a simple feed-forward MLP with 2 hidden layers. It takes as input the concatenation of the node embeddings of the first and last nodes on the current route, $\mathbf{s}_{t}$ and $\tau_{r_{t}}$. It outputs a scalar $h$. We apply the sigmoid function to $h$ to get the halting policy:
@@ A.2.3 Extension module
The extension module $\textup{NN}_{ext}$ computes the node-pair scores $o_{ij}$ that are used to calculate the probabilities of selecting each extension on odd-numbered timesteps.
This should be a function of the time between nodes along the path, $\tau_{aij}$, as well as the demand between $i$ and $j$, and potentially of other features of $\mathcal{C}$ and $G_{\mathcal{R}}$. But since the paths are all shortest paths, $\tau_{aij}=T_{ij}$ in every case, so the score $o_{aij}$ gained by linking nodes $i,j$ is independent of the chosen path $a$: $o_{aij}=o_{ij}$. Let these be computed by some function $f(\cdot)$:
$o_{ij}=\begin{cases}f(T_{ij},i,j,\mathcal{C},G_{\mathcal{R}}),&\text{if}i\neq j\\ 0,&\text{if}i=j\end{cases}$ (16)
Assuming these node-pair scores are an accurate indication of the benefit of linking those nodes, the score of the path as a whole should simply be the sum of scores of the node-pairs it links:
$o_{a}=\sum_{k\in a}\sum_{l\in a}o_{kl}$ (17)
Then, suppose we have a partially-planned route $r$ and are considering paths $a\in\textup{SP}$ to extend $r$. Each of these paths still links all the nodes along it, so can achieve score $o_{a}$. But appending $a$ to $r$ will also link every node in $r$ to every node in $a$, so the benefit of extending $r$ by $a$ should depend on these as well. However, the time taken to get from node $i\in r$ to node $j\in a$ may be greater than $T_{ij}$, since $r|a$ is not necessarily a shortest path. So the shortest-path node-pair scores $o_{ij}$ here are not appropriate. Instead, new node-pair scores $o_{(r|a)ij}$ should be computed based on the driving time $\tau_{(r|a)ij}$ that would result from joining $a$ to $r$. Otherwise, the same function $f(\cdot)$ should be appropriate, with the new inter-node time:
$o_{(r|a)ij}=\begin{cases}f(\tau_{(r|a)ij},i,j,\mathcal{C},G_{\mathcal{R}}),&\text{if}i\neq j\\ 0,&\text{if}i=j\end{cases}$ (18)
And the overall score should then be:
$o_{(r|a)}=o_{a}+\sum_{i\in r}\sum_{j\in a}\Delta_{i\neq j}o_{(r|a)ij}$ (19)
Where $\Delta_{i\neq j}$ is 0 if $i=j$ and 1 otherwise, and serves to avoid including the score for the overlapping end-node of $r$ and $a$ in the sum.
We design the extension policy head to compute the score function $f(\cdot)$. In this policy head, the function $f(\cdot)$ that computes node-pair scores is an MLP, called $\text{MLP}_{ext1}$, that takes as input $T_{ij}$ or $\tau_{(r|a)ij}$ (as appropriate), the backbone GAT’s embeddings $y_{i}$ and $y_{j}$ of nodes $i$ and $j$, and the global state vector $\mathbf{s}$. For each path $a\in\textup{SP}$, we sum these node-pair scores to get path scores $o_{a}$ as in Equation 17, and when extending a route $r$, we compute the extension scores $o_{(r|a)}$ as in Equation 19. We can efficiently compute the node-pair scores in a batch by pairing all node embeddings $y_{i}$ when $r_{t}=[]$; and when $r_{t}\neq[]$, we only compute scores $o_{(r|a)ij}$ for node pairs where exactly one of $i$ or $j$ is in $r_{t}$. The resultant scalar scores, $o_{a}$ and $o_{(r|a)}$, are provided to the next stage of the extension policy head.
@@
If $\alpha=0$, then the benefit of choosing a path $a$ to start or extend a route depends on its driving time $\tau_{a}$, and generally on the state $s$. And if $10<\alpha<$ 1, then the benefit of $a$ depends on these as well as on the edges it adds to $\mathcal{E}_{\mathcal{R}}$, and thus on $o_{(r|a)}$. So our architecture ought to incorporate all of these factors. We achieve this with a final MLP, called $\text{MLP}_{ext2}$, that takes as input $\mathbf{s}_{t},\tau_{a},$ and the previous node-pair-based score ($o_{a}$ or $o_{(r|a)}$, depending on whether $r_{t}=[]$), and outputs a final scalar score $\hat{o}_{a}$ for each candidate path. The extension policy is then the softmax of these values:
$\pi_{\theta}(a|s)=\frac{e^{\hat{o}_{a}}}{\sum_{a^{\prime}\in\mathcal{A}}e^{\hat{o}_{a^{\prime}}}}$ (20)
We thus treat the outputs of $\textup{NN}_{ext}$ as un-normalized log-probabilities of the possible actions during an extension step.
A.3 Value Network